### v5

# Towards a Theory of Pragmatic Information

Abstract
Standard information theory says nothing about how much meaning is conveyed by a message. We fill this gap with a rigorously justifiable, quantitative definition of “pragmatic information”, the amount of meaning in a message relevant to a particular decision. We posit that such a message updates a random variable, $\omega$, that informs the decision. The pragmatic information of a single message is then defined as the Kulbach-Leibler divergence between the apriori and aposteriori probabilities of $\omega$; the pragmatic information of a message ensemble is the expected value of the pragmatic information of the ensemble’s component messages. We justify these definitions by proving that the pragmatic information of a single message is the expected difference between the shortest binary encoding of $\omega$ under the a priori and a posteriori distributions, and that the average of the pragmatic values of individual messages, when sampled a large number of times from the ensemble, approaches its expected value.
Pragmatic information is non-negative and additive for independent decisions and “pragmatically independent” messages. Also, pragmatic information is the information analogue of free energy: just as free energy quantifies the part of a system’s total energy available to do useful work, so pragmatic information quantifies the information actually used in making a decision.
We sketch 3 applications: the single play of a slot machine, a.k.a. a “one armed bandit”, with an unknown payout probability; a characterization of the rate of biological evolution in the so-called “quasi-species” model; and a reformulation of the efficient market hypothesis of finance. We note the importance of the computational capacity of the receiver in each case.
1 Introduction
Just about any discussion of standard information theory notes that the standard entropy measure of the amount of information in a transmitted message says nothing about how meaningful the message is, even though a non-technical person would say that a message is “informative” if they are able to extract meaning from it. Warren Weaver’s introduction to Shannon’s foundational paper [2] addresses this point by observing that the effectiveness of a communications process could be measured by answering any of the following three questions:
  1. A.
    How accurately can the symbols that encode the message be transmitted (“the technical problem”)?
  2. B.
    How precisely do the transmitted symbols convey the desired meaning (“the semantics problem”)?
  3. C.
    How effective is the received message in changing conduct (“the effectiveness problem”)?
Weaver then observes that Shannon’s paper — and thus the entire edifice of what is now known as “information theory” — concerns itself only with the answer to above question A [^1] Arguably, however,every answer to question A is implicitly an answer to question C, as every transmission implies that something in the receiver, which might be called its “conduct”, must have changed.. The present paper, in contrast, is a continuation of previous work [3], in which it was proposed that the only definitive way to assess the “meaning” of question B above is to quantify the “changed conduct” of question C [^2] Such a quantification begs the question of whether this change in conduct is actually a change for the better! Indeed, recent events have shown that dis-information can be quite effective in changing conduct, an observation that will be taken into account below..
Although the primary goal of the present work is to advance the theory as such, any theory that is termed “pragmatic” should have practical applications, so we mention a few. [3] proposed that the rate at which biological evolution aggregates “meaningful information” could be used to measure the rate at which evolution proceeds. Throughout the present paper, we present a second application, namely the role that the history of previous payouts of a slot machine acquire meaning in the context of deciding whether to keep playing. We also show how our conceptual framework makes possible a reformulation of the celebrated efficient market hypothesis of financial economics as the claim that none of the available information about a given security is “pragmatically useful”.
In the next section, we discuss the pragmatic information of a single message, preparing our quantitative results with a discussion of the properties that pragmatic information should have. Some of these properties have been noted by other authors, qualitatively by such authors as [6] and [7] and more quantitatively by [8] and [9], the latter having proposed definitions of pragmatic information similar to ours. However, they do not provide anything like our demonstration that this definition follows from natural assumptions about what a “measure of meaning” should be.
We then consider the pragmatic information of messages in an ensemble. As part of that discussion, we provide an example of how our definition applies to the simplest multi-armed bandit problem, that of playing a single slot machine with an unknown payout probability. We also consider what it means for messages to be “pragmatically independent” and the relationship between the pragmatic information of an ensemble of messages and the mutual information between the message ensemble and the “state of the world” variable $\omega$.
We conclude with speculation about how a kind of “computational capacity” may play a role in the theory of pragmatic information similar to that of channel capacity in the Shannon theory, in that beliefs about $\omega$ cannot be updated during the non-zero time required to complete message processing. Given that a participant in the financial markets can only find market information useful if it is meaningful in our sense, we also discuss our ideas in the context of the efficient market hypothesis.
2 The Pragmatic Information of a Single Message
2.1 Our Conceptual Framework
As with the Shannon theory of information, we define a message to be a finite symbol string in some alphabet, albeit with the additional implication that meaning is being communicated. If the practical meaning of such a message stems from its usefulness in making a decision, the message must do so by changing those beliefs about the state of the world that will inform the subsequent action. Accordingly, we consider a decision maker, $\Delta$, which may be a machine, a human, or a non-human organism. $\Delta$’s decision is informed by its limited knowledge of the state of the world. We represent this limited knowledge as the probability distribution of a random variable $\omega$ in some discrete sample space $\Omega=\{\omega_{1},\omega_{2},\ldots,\omega_{i},\ldots,\omega_{N}\}$. Initially, $\Delta$ assigns a priori probabilities $\mathbf{q}=(q_{1},q_{2},\ldots,q_{N})$, each of which are positive, to the various possible values of $\omega$. We refer to $\mathbf{q}$ as a priori probabilities because we then supply $\Delta$ with a “message”, $m$. As a result, $\Delta$’s beliefs about the state of the world change from $\mathbf{q}$ to $\mathbf{p}_{m}=(p_{1|m},p_{2|m},\ldots,p_{N|m})$.
While our conceptual framework can be applied to the wealth of so-called “multi-armed bandit problems” (See, for example, [11] and [12].), we consider the following, even simpler “one-armed bandit” problem: we can pay \$1 to play a slot machine (also known as a “one armed bandit”) as many times as we wish. After each play, we will win a \$2 payout with some constant probability $\pi$, or nothing, with probability $1-\pi$, with the payout events independent from one play to the next. The problem is that $\pi$ is not known in advance!
Suppose we denote the result of the $t^{th}$ such play by the random variable $\mathbf{1}_{t}$, with
$\mathbf{1}_{t}=\begin{cases}1&\quad\text{if there is a payout on the $t^{th}$ play}\\ 0&\quad\text{otherwise}\end{cases}$
Clearly, we learn something about $\pi$, and thus the correct play/no play decision for subsequent trials from the realized values of the sequence $\mathbf{1}_{1},\mathbf{1}_{2},\ldots,\mathbf{1}_{T}$ , but exactly how much? If we have won $w$ times in the $T$ trials, Laplace’s Rule of Succession [14] estimates $\pi$ as
$\hat{\pi}(w,T)=\frac{w+1}{T+2}.$ (1)
$\hat{\pi}(w,T)$ is less and less sensitive to whether or not any individual trial results in a payout as $T$ increases. Thus, we learn less and less about what we need to know from each additional value in this sequence; indeed, after a finite number of such values, we know essentially all we need to know to make our decision.
In contrast, the amount of Shannon information,
$\mathcal{H}(\mathbf{1}_{t})=\pi\log_{2}\pi+(1-p)\log_{2}(1-\pi),$
per play remains constant from one play to the next, so that the Shannon information for the $T$ independent plays is $T[\pi\log_{2}\pi+(1-\pi)\log_{2}(1-\pi)]$. However, the Shannon information captures data such as the particular order of successes and failures, which is more than we need to know to estimate $\pi$.
2.2 Defining the Pragmatic Information of a Single Message
Informed by the above, we propose the following
Definition.
The pragmatic information, $D_{\Delta}(\mathbf{p}_{m}||\mathbf{q})$, of a single message $m$ acting on $\Delta$ is [^3] As is customary in writing about information theory, we assume base 2 logarithms here and subsequently.
$D_{\Delta}(\mathbf{p}_{m}||\mathbf{q})=\sum_{i}p_{i|m}\log\left(\frac{p_{i|m}}{q_{i}}\right),$ (2)
with the usual convention that $p_{i|m}\log\left(\frac{p_{i|m}}{q_{i}}\right)=0$ if $p_{i|m}=0$. In other words, $D_{\Delta}(\mathbf{p}_{m}||\mathbf{q})$ is the Kullbach-Liebler divergence of $\mathbf{p}_{m}$ and $\mathbf{q}$.
As we will see in more detail below, our definition of pragmatic information incorporates many of the features postulated by [6] and [7], such as:
  • $D_{\Delta}(\mathbf{p}_{m}||\mathbf{q})=0$ if $m$ is already known to $\Delta$.
  • $D_{\Delta}(\mathbf{p}_{m}||\mathbf{q})=0$ if if $\Delta$ cannot process $m$.
  • $D_{\Delta}(\mathbf{p}_{m}||\mathbf{q})$ will depend on $\Delta$, and, more specificially, on the context in which $\Delta$ receives $m$.
  • For two different messages, $m$ and $m^{\prime}$,
    $D_{\Delta}(\mathbf{p}_{m}||\mathbf{q})=D_{\Delta}(\mathbf{p}_{m^{\prime}}||\mathbf{q})$
    if $m$ and $m^{\prime}$ always lead $\Delta$ to choose the same decision probabilities, even if $m$ and $m^{\prime}$ have different lengths and/or content.
  • $D_{\Delta}(\mathbf{p}_{m}||\mathbf{q})$ increases if $\Delta$ is able to recognize the increasing novelty of $m$ as $\mathbf{p}_{m}$ differs more from $\mathbf{q}$. However, if $m$ is completely novel, in the sense that $\Delta$ will be unable to process it, $D_{\Delta}(\mathbf{p}_{m}||\mathbf{q})=0$. This situation is considered in more detail below.
Our conceptual framework also allows us to look more carefully at the following example, introduced by [7]:
Consider a cash register which stores the data of all sales over the day. After closing time, the shopkeeper is not interested in the list of numbers, but in the total. By performing the addition the cash register produces pragmatic information, since that total comes closer to the user requirements than the list of raw data would do.
In our framework, $\Delta$ would be the shopkeeper who, presumably, needs to make decisions based on the true sales volume for the day, which we take as the relevant random variable $\omega$. $\Delta$ would have an a priori estimate of $\omega$, based on the list of individual sales that is reflected in an a priori probability distribution, $\mathbf{q}$. This distribution, when updated upon learning the cash register output of total sales, $S_{T}$, for the day, is likely to become more peaked around a single value. Thus, if $\mathbf{p}_{T}$ is this updated distribution, $D_{\Delta}(\mathbf{p}_{T}||\mathbf{q})>0$.
However, everything depends on the shopkeeper’s subjective beliefs! The shopkeeper might believe that the buggy software of the cash register is less reliable than the shopkeeper’s own manual addition, in which case $S_{T}$ will be ignored and $D_{\Delta}(\mathbf{p}_{T}||\mathbf{q})=0$.
The formula that [9] proposed for “surprise” is essentially the same as (2) for a single message (Below, we will generalize this definition to any $m$ in some ensemble of possible messages.). However, [9] left ambiguous whether to define this quantity as (in our notation) $D_{\Delta}(\mathbf{p}_{m}||\mathbf{q})$ or $D_{\Delta}(\mathbf{q}||\mathbf{p}_{m})$. They prefer the latter, as it easier for them to compute intuition-building analytic results in specific cases; however, we prefer the former, as we assume – perhaps naively! – that $\mathbf{p}_{m}$ really does provide an improvement over $\mathbf{q}$. Our preferred definition also makes more sense in subsequent sections of this paper, where we list some known/easily derivable properties [^4] See [15] for proofs of all of the theorems in the next subsection, except for our “Wrong Code Theorem”, which is an embellishment of their results, and the obvious corollary to Theorem 2.3.4. of $D_{\Delta}(\mathbf{p}_{m}||\mathbf{q})$ that make it an appealing choice for a measure of pragmatic information of a single message.
2.3 Properties of the Pragmatic Information of a Single Message
The fundamental justification for our definition of the pragmatic information of a single message is found in the following two theorems and a corollary to the second:
Theorem 2.3.1 (Wrong Code Theorem).
Suppose $L_{\mathbf{q}}(\omega)$ is the length, in bits, of the shortest binary code required to communicate that $\Delta$ has decided upon outcome $\omega$, assuming the prior probabilities $\mathbf{q}$, and suppose $L_{\mathbf{p}_{m}}(\omega)$ is the corresponding length, assuming the prior probabilities $\mathbf{p}_{m}$. Then
${\mathcal{E}}\left[L_{\mathbf{q}}(\omega)-L_{\mathbf{p}_{m}}(\omega)\right]=D_{\Delta}(\mathbf{p}_{m}||\mathbf{q}),$
where ${\mathcal{E}}\left[L_{\mathbf{q}}(\omega)-L_{\mathbf{p}_{m}}(\omega)\right]$ is the expected length of $L_{\mathbf{q}}(\omega)-L_{\mathbf{p}_{m}}(\omega)$ under the a posteriori probabilities $\mathbf{p}_{m}$.
Proof.
Let ${\mathcal{H}}(\mathbf{p}_{m})$ be the Shannon entropy of $\mathbf{p}_{m}$. Then, per [15],
${\mathcal{H}}(\mathbf{p}_{m})+D(\mathbf{p}_{m}||\mathbf{q})\leq{\mathcal{E}}\left[L_{\mathbf{q}}(\omega)\right]]<{\mathcal{H}}(\mathbf{p}_{m})+D_{\Delta}(\mathbf{p}_{m}||\mathbf{q})+1,$
and
${\mathcal{H}}(\mathbf{p}_{m})\leq{\mathcal{E}}\left[L_{\mathbf{p}_{m}}(\omega)\right]<{\mathcal{H}}(\mathbf{p}_{m})+1,$
The desired result follows upon subtracting the second set of inequalities from the first. ∎
We conclude that $D_{\Delta}(\mathbf{p}_{m}||\mathbf{q})$ can reasonably be interpreted as the amount of information that $\Delta$ has “learned” from $m$.
Next, consider two decision makers, $\Delta$ and $\Delta^{\prime}$, with respective input messages $m$ and $m^{\prime}$, estimating the likelihood of random variables $\omega\in\Omega$ and $\omega^{\prime}\in\Omega^{\prime}$, respectively. Suppose:
  • $\{q_{i,i^{\prime}}\}$ is the set ofa priori joint probabilities regarding outcome $\omega_{i}\in\Omega$ and, simultaneously, outcome $\omega^{\prime}_{i^{\prime}}\in\Omega^{\prime}$, each without knowledge of respective input messages $m$ and $m^{\prime}$.
  • the action of both messages changes $\{q_{i,i^{\prime}}\}$ to the a posteriori joint probabilities, $\{p_{i,i^{\prime}|m,m^{\prime}}\}$.
  • $\{q^{\prime}_{i^{\prime}|i}\}$ is the set of prior probabilities of $\Delta^{\prime}$’s decisions, given the decisions of $\Delta$, but not message $m$, and $\{p^{\prime}_{i^{\prime}|i,m,m^{\prime}}\}$ is the a posteriori probability distribution of $\Delta^{\prime}$’s decisions, given the decisions of $\Delta$ and both messages $m$ and $m^{\prime}$.
We then have
Theorem 2.3.2 (Chain Rule for Kullbach-Leibler Divergence).
$D(\{p_{i,i^{\prime}|m,m^{\prime}}\}||\{q_{i,i^{\prime}}\})=D(\mathbf{p}_{m}||\mathbf{q})+D(\{p^{\prime}_{i^{\prime}|i,m,m^{\prime}}\}||\{q^{\prime}_{i^{\prime}|i}\})$
Corollary (Additivity of Kullbach-Leibler Divergence).
In the special case where $\Delta$ and $\Delta^{\prime}$ are completely independent of each other, i.e. that $\{q^{\prime}_{i^{\prime}|i}\}=\mathbf{q}^{\prime}$ and $\{p^{\prime}_{i^{\prime}|i,m,m^{\prime}}\}=\mathbf{p}^{\prime}_{m}$,
$D(\{p_{i,i^{\prime}|m,m^{\prime}}\}||\{q_{i,i^{\prime}}\})=D(\mathbf{p}_{m}||\mathbf{q})+D(\mathbf{p}^{\prime}_{m^{\prime}}||\mathbf{q}^{\prime}),$
about which Kullbach writes in [16] that
“Additivity of information for independent events is intuitively a fundamental requirement, and is indeed postulated in most axiomatic developments of information theory. Additivity is the basis for the logarithmic form of information. A sample of $n$ independent observations from the same population [here $n$ copies of $\Delta$ making $n$ identical and independent decisions] provides $n$ times the mean information in a single observation.”
$D_{\Delta}(\mathbf{p}_{m}||\mathbf{q})$ has a few other properties that would seem to be required of a measure of meaning, as stated in the following theorems:
Theorem 2.3.3 (Non-negativity of Kullbach-Leibler Divergence).
$D_{\Delta}(\mathbf{p}_{m}||\mathbf{q})\geq 0,$
with equality if and only if $\mathbf{p}_{m}=\mathbf{q}$.
Thus, $m$ always conveys positive pragmatic information, unless $m$ is ignored by $\Delta$.
Theorem 2.3.4 (Convexity of Kullbach-Leibler Divergence).
Consider the pairs of probability distributions $\mathbf{p}_{m}$ and $\mathbf{p}^{\prime}_{m}$ and $\mathbf{q}$, and $\mathbf{q}^{\prime}$. Form, for any $\lambda\in[0,1]$, the pair of interpolated distributions
$\mathbf{p}_{m}^{\lambda}=\lambda\mathbf{p}_{m}+(1-\lambda)\mathbf{p}^{\prime}_{m}$
and
$\mathbf{q}^{\lambda}=\lambda\mathbf{q}+(1-\lambda)\mathbf{q}^{\prime},$
then
$D\Big{(}\mathbf{p}_{m}^{\lambda}||\mathbf{q}^{\lambda}\Big{)}\leq\lambda D\left(\mathbf{p}_{m}||\mathbf{q}\right)+(1-\lambda)D\left(\mathbf{p}^{\prime}_{m}||\mathbf{q}^{\prime}\right).$
In the special case where $\mathbf{q}=\mathbf{q}^{\prime}=\mathbf{q}_{\lambda}$ and ${\mathbf{u}}_{j}$ and ${\mathbf{u}}_{k}$ are unit vectors in the $j^{\rm th}$ and $k^{\rm th}$ direction, respectively,
$D\Big{(}\lambda{\mathbf{u}}_{j}+(1-\lambda){\mathbf{u}}_{k}||\mathbf{q}\Big{)}$ $\leq\lambda D\left({\mathbf{u}}_{j}||\mathbf{q}\right)+(1-\lambda)D\left({\mathbf{u}}_{k}||\mathbf{q}\right)$
$\leq\max\left(-\log q_{j},-\log q_{k}\right).$
It follows that, for given $\mathbf{q}$, we have the following
Corollary.
The maximum value of $D_{\Delta}(\mathbf{p}_{m}||\mathbf{q})$ is $\max_{l}(-\log q_{l})$, with the max taken over all of the $q$’s for which $p_{l|m}>0$.
In other words, the pragmatic information conveyed by a single message $m$ is a maximum when $\Delta$ treats the outcome indicated by $m$ as a certainty.
3 The Pragmatic Information of an Ensemble of Messages
We now assume that $\Delta$ is asked to make a series of independent decisions, each time given the same prior set of beliefs (i.e. $\mathbf{q}$ is the same each time), but each time given a (possibly different) message $M_{k}$ from some finite ensemble $\mathcal{M}$. We further assume that each $M_{k}$ is sampled using the same probability distribution, namely ${\rm Pr}\left\{M_{k}=m\right\}=\varphi_{m}$ for all $k$ and every $m\in\mathcal{M}$ is eventually sampled, regardless of how the sampling starts. Given the additivity of Kullbach-Leibler divergence for independent decisions, $M_{k}$ conveys the (possibly different) amount of pragmatic information $D_{\Delta}(\mathbf{p}_{M_{k}}||\mathbf{q})$ for each $k$. The average amount of pragmatic information that $\Delta$ accrues per message is then the random variable
$\Phi_{N}={1\over N}\sum_{k=1}^{N}D_{\Delta}(\mathbf{p}_{M_{k}}||\mathbf{q}).$
Our assumptions on how the $M_{k}$’s are sampled are sufficient to guarantee that this sum satisfies Birkoff’s ergodic theorem [10]. We can therefore conclude that
$\lim_{N\rightarrow\infty}\Phi_{N}=E\left[D(\mathbf{p}_{M}||\mathbf{q})\right],$
with probability 1, with the subscript being omitted from the expectation because the distribution of the $M_{k}$’s is the same for all $k$. We stress that, as in the previous section, $\Delta$ makes each of its decisions as though none of the previous decisions had ever been made. However, the hypotheses of the ergodic theorem are sufficiently weak that no such assumption is required for the $M_{k}$’s. While they could be sampled independently, they could also be a realization of a Markov chain, etc. The generality of this result therefore justifies the following
Definition.
The pragmatic information, ${\Phi_{\Delta}(\mathcal{M};\Omega)}$, of an ensemble of messages ${\mathcal{M}}$ is the expected value
$\Phi_{\Delta}(\mathcal{M};\Omega)$ $=E\left[D(\mathbf{p}_{m}||\mathbf{q})\right]$
$=\sum_{i,m}\varphi_{m}p_{i|m}\log\left(\frac{p_{i|m}}{q_{i}}\right)$
$=\sum_{i,m}p_{i,m}\log\left(\frac{p_{i|m}}{q_{i}}\right)$
where $p_{i|m}$ remains the conditional probablilty of outcome $\omega_{i}$, given the receipt of message $m$, so that the joint probability, $p_{i,m}$, of outcome $\omega_{i}$ upon receiving message $m$ is $p_{i,m}=\varphi_{m}p_{i|m}$. Once again, we take $p_{i|m}\log\left(\frac{p_{i|m}}{q_{i}}\right)=0$ if $p_{i|m}=0$.
3.1 An Example: The One Armed Bandit Problem
It is instructive to apply this definition to the “one armed bandit problem” discussed above. Evidently, the relevant decision is whether it is worth playing the slot machine, a decision that is informed by a random variable, $\omega$, that can assume exactly one of the two values $\mathtt{PAYOUT}$ and $\mathtt{NOPAYOUT}$, i.e. $\Omega=\{\mathtt{PAYOUT},\mathtt{NOPAYOUT}\}$. We want to know the pragmatic information accruing from a knowledge of $\mathbf{1}_{T+1}$, given that the $T$ prior trials have resulted in $w$ wins. Thus, per (1), we take $\hat{\pi}(w,T)$ as $q_{1}$, our a priori estimate of the payout probability, and $1-\hat{\pi}(w,T)$ as $q_{0}$, our a priori estimate of the no payout probability.
The ensemble, $\mathcal{M}_{T+1}$, of possible messages for the $T+1^{st}$ trial is $\mathtt{PAYOUT}_{T+1}$ and $\mathtt{NOPAYOUT}_{T+1}$. If $\Delta$ receives the $\mathtt{PAYOUT}_{T+1}$ message, the a posteriori $\mathtt{PAYOUT}$ and $\mathtt{NOPAYOUT}$ probabilities are updated to
$\mathbf{\pi}_{1}$ $=\Big{(}\hat{\pi}(w+1,T+1),1-\hat{\pi}(w+1,T+1)\Big{)}$
$=\Big{(}\frac{w+2}{T+3},\ \frac{T-w+1}{T+3}\Big{)}$
Similarly, if $\Delta$ receives the $\mathtt{NOPAYOUT}_{T+1}$ message, the corresponding a posteriori probability vector is
$\mathbf{\pi}_{0}$ $=\Big{(}\hat{\pi}(w+1,T+1),1-\hat{\pi}(w+1,T+1)\Big{)}$
$=\Big{(}\frac{w+1}{T+3},\ \frac{T-w+2}{T+3}\Big{)}$
We then have
$D_{\Delta}(\mathbf{\pi}_{1}||\mathbf{q})$ $=\hat{\pi}(w+1,T+1)\log\left[\frac{\hat{\pi}(w+1,T+1)}{\hat{\pi}(w,T)}\right]+$
$\qquad\left[1-\hat{\pi}(w+1,T+1)\right]\log\left[\frac{1-\hat{\pi}(w+1,T+1)}{1-\hat{\pi}(w,T)}\right]$
$=\frac{w+2}{T+3}\log\left[\frac{(w+2)(T+2)}{(T+3)(w+1)}\right]+\frac{T-w+1}{T+3}\log\left[\frac{T+2}{T+3}\right]$
and
$D_{\Delta}(\mathbf{\pi}_{0}||\mathbf{q})$ $=\hat{\pi}(w,T+1)\log\left[\frac{\hat{\pi}(w,T+1)}{\hat{\pi}(w,T)}\right]+$
$\qquad\left[1-\hat{\pi}(w,T+1)\right]\log\left[\frac{1-\hat{\pi}(w,T+1)}{1-\hat{\pi}(w,T)}\right]$
$=\frac{w+1}{T+3}\log\left[\frac{T+2}{T+3}\right]+\frac{T-w+2}{T+3}\log\left[\frac{(T-w+2)(T+2)}{(T+3)(T-w+1)}\right],$
both of which can be interpreted as the reduction in the uncertainty surrounding the estimate of $\pi$, upon receipt of the corresponding value of $\mathbf{1}_{T+1}$. We conclude that the pragmatic information provided by the $T+1^{st}$ trial is
$\Phi_{\Delta}\left(\mathcal{M}_{T+1};\Omega\right)$ $=\pi D_{\Delta}(\mathbf{\pi}_{1}||\mathbf{q})+(1-\pi)D_{\Delta}(\mathbf{\pi}_{0}||\mathbf{q})$
$=p\left\{\frac{w+2}{T+3}\log\left[\frac{(w+2)(T+2)}{(T+3)(w+1)}\right]+\frac{T-w+1}{T+3}\log\left[\frac{T+2}{T+3}\right]\right\}$
$\qquad+(1-\pi)\left\{\frac{w+1}{T+3}\log\left[\frac{T+2}{T+3}\right]+\frac{T-w+2}{T+3}\log\left[\frac{(T-w+2)(T+2)}{(T+3)(T-w+1)}\right]\right\}$
As $T$ gets large, $w=\pi T+o(T)$ with probability 1, by the Kolmogorov Strong Law of Large Numbers [17]. Straightforward algebraic manipulation then shows that the arguments of all of the logarithms in $\Phi_{\Delta}\left(\mathcal{M}_{T+1};\Omega\right)$ approach 1 monotonically as $T$ increases. We conclude that $\Phi_{\Delta}\left(\mathcal{M}_{T+1};\Omega\right)$ steadily decreases to zero with increasing $T$.
3.2 Properties of the Pragmatic Information of an Ensemble of Messages
The theorems in the previous section remain true for ${\Phi_{\Delta}(\mathcal{M};\Omega)}$, since it is just a weighted sum of Kullbach-Leibler divergences. We therefore have immediately
Theorem 3.2.1 (Non-negativity of Pragmatic Information).
$\Phi_{\Delta}(\mathcal{M};\Omega)\geq 0$, with equality if and only if none of the messages in $\mathcal{M}$ change $\Delta$’s a priori estimate of $\omega$.
Theorem 3.2.2 (Upper Bound on Pragmatic Information).
$\Phi_{\Delta}(\mathcal{M};\Omega)$ $\leq-\sum_{m}\varphi_{m}\log q_{k(m)}$
$\leq-\log q_{l}$
where, for each $m$, $k(m)=\operatorname*{arg\,max}(-\log q_{l})$, with the max taken over all of the $q$’s for which $p_{l|m}>0$ and $-\log q_{l}$ is the largest of the $-\log q$’s (The $\varphi$’s sum to 1.).
and
Theorem 3.2.3 (Wrong Code Theorem for Pragmatic Information).
$\Phi_{\Delta}(\mathcal{M};\Omega)$ is the expected number of extra bits required to code samples from $\mathbf{p}_{m}$, averaged over $m\in\mathcal{M}$, using a binary encoding optimized for $\bf q$, rather than, for each $m$, a binary encoding optimized for that $\mathbf{p}_{m}$.
Next, we generalize joint and conditional Kullbach-Leibler Divergence. Suppose:
  • decision makers $\Delta$ and $\Delta^{\prime}$ have respective “state of the world” variables $\omega\in\Omega$ and $\omega^{\prime}\in\Omega^{\prime}$,
  • $\{q_{i,i^{\prime}}\}$ is the set of a priori joint probabilities that $\Delta$ assigns to $\omega_{i}\in\Omega$ and $\Delta^{\prime}$ assigns to $\omega^{\prime}_{i^{\prime}}\in\Omega^{\prime}$, prior to receipt of respective input messages $m$ and $m^{\prime}$,
  • the action of both messages changes the joint probabilities of outcomes $\omega_{i}$ and $\omega^{\prime}_{i^{\prime}}$ from the prior probabilities, $\{q_{i,i^{\prime}}\}$, to the a posteriori joint probabilities, $\{p_{i,i^{\prime}|m,m^{\prime}}\}$, and
  • $\{q^{\prime}_{i^{\prime}|i}\}$ is the set of prior probabilities of $\Delta^{\prime}$’s decisions, given the decisions of $\Delta$, but not message $m$, and $\{p^{\prime}_{i^{\prime}|i,m,m^{\prime}}\}$ is the a posteriori probability distribution of $\Delta^{\prime}$’s decisions, given the decisions of $\Delta$ and both messages $m$ and $m^{\prime}$.
  • The joint probability that $\Delta$ receives message $m$ and $\Delta^{\prime}$ receives message $m^{\prime}$ is $\varphi_{m,m^{\prime}}$.
  • The joint probability that these messages are received and the decision makers output $\omega_{i}$ and $\omega^{\prime}_{i^{\prime}}$, respectively, is $p_{i,i^{\prime},m,m^{\prime}}$.
We then have the following
Definition.
The joint pragmatic information of messages in ${\mathcal{M}}$ acting on $\Delta$ and messages in ${\mathcal{M}^{\prime}}$ acting on $\Delta^{\prime}$ is
$\Phi_{\Delta,\Delta^{\prime}}(\mathcal{M},\mathcal{M}^{\prime};\Omega,\Omega^{\prime})$ $=E\left[D(\{p_{i,i^{\prime}|M,M^{\prime}}\}||\{q_{i,i^{\prime}}\})\right]$
$=\sum_{i,i^{\prime},m,m^{\prime}}\varphi_{m,m^{\prime}}p_{i,i^{\prime}|m,m^{\prime}}\log\left(\frac{p_{i,i^{\prime}|m,m^{\prime}}}{q_{i,i^{\prime}}}\right)$
$=\sum_{i,i^{\prime},m,m^{\prime}}p_{i,i^{\prime},m,m^{\prime}}\log\left(\frac{p_{i,i^{\prime}|m,m^{\prime}}}{q_{i,i^{\prime}}}\right)$
and
Definition.
The conditional pragmatic information of messages in ${\mathcal{M}^{\prime}}$ acting upon $\Delta^{\prime}$, given that $m\in\mathcal{M}$ acting on $\Delta$, as well as the resulting outcome $\omega$ is
$\Phi_{\Delta^{\prime}|\Delta}(\mathcal{M}^{\prime};\Omega^{\prime}|\mathcal{M},\Omega)$ $=E\left[D(\{p_{i^{\prime}|i,M,M^{\prime}}\}||\{q_{i^{\prime}|i}\})\right]$
$=\sum_{i,i^{\prime},m,m^{\prime}}p_{i,i^{\prime},m,m^{\prime}}\log\left(\frac{p^{\prime}_{i^{\prime}|i,m,m^{\prime}}}{q_{i^{\prime}|i}}\right).$
Since the Kullbach-Leibler divergence of each $m\in\mathcal{M}$ satisfies the Chain Rule for Kullbach-Leibler divergence, the generalization of that Rule follows immediately:
Theorem 3.2.4 (Chain Rule for Pragmatic Information).
Suppose each $m\in\mathcal{M}$ and $m^{\prime}\in\mathcal{M}^{\prime}$ acts, respectively, on decision makers $\Delta$ and $\Delta^{\prime}$, with respective prior decision probabilities $\mathbf{q}$ and $\mathbf{q}^{\prime}$ and with corresponding a posteriori decision probabilities $\mathbf{p}_{m}$ and $\mathbf{p}^{\prime}_{m^{\prime}}$, then
$\Phi_{\Delta,\Delta^{\prime}}(\mathcal{M};\Omega,\Omega^{\prime})=\Phi_{\Delta}({\mathcal{M}};\Omega)+\Phi_{\Delta^{\prime}|\Delta}(\mathcal{M}^{\prime};\Omega^{\prime}|\mathcal{M},\Omega)$
As an example, suppose $\Delta$ and $\Delta^{\prime}$ have the respective Boolean state of the world variables $\omega$ and $\omega^{\prime}$, informed by Boolean messages $M$ and $M^{\prime}$ according to the formulas $\omega=M$ and $\omega^{\prime}=M\ \mathtt{AND}\ M^{\prime}$ with probability one. Assume further that $M$ and $M^{\prime}$ are independent random variables, each assuming the values 1 and 0 with equal probability. Thus, a priori, $\omega$ also assumes the values 1 and 0 with equal probability, $\omega^{\prime}$ is 1 with probability ${1\over 4}$ and 0 with probability ${3\over 4}$, and $p_{i,i^{\prime}|m,m^{\prime}}=p_{i^{\prime}|i,m,m^{\prime}}=1$ for $(i,i^{\prime},m,m^{\prime})=(0,0,0,0)$, $(i,i^{\prime},m,m^{\prime})=(0,0,0,1)$, $(i,i^{\prime},m,m^{\prime})=(1,0,1,0)$, and $(i,i^{\prime},m,m^{\prime})=(1,1,1,1)$ and zero otherwise. We then have
$\Phi_{\Delta,\Delta^{\prime}}(\mathcal{M};\Omega,\Omega^{\prime})={3\over 2},$
with
$\Phi_{\Delta}({\mathcal{M}};\Omega)=1,$
and
$\Phi_{\Delta^{\prime}|\Delta}(\mathcal{M}^{\prime};\Omega^{\prime}|\mathcal{M},\Omega)={1\over 2}.$
We also note that conditioning need not decrease pragmatic information, as it does with Shannon entropy, since
$\Phi_{\Delta^{\prime}}({\mathcal{M}^{\prime}};\Omega^{\prime})$ $={3\over 2}-{3\over 4}\log 3\approx 0.3113$
$<{1\over 2}=\Phi_{\Delta^{\prime}|\Delta}(\mathcal{M}^{\prime};\Omega^{\prime}|\mathcal{M},\Omega)$
In this case, the receipt of $M$ has provided a context in which $M^{\prime}$ has become more informative.
3.3 Pragmatic Independence of Messages
It is well known that the joint Shannon entropy of two random variables is the sum of their individual entropies if and only if the random variables are probabalistically independent. Here, we consider the corresponding situation for pragmatic information, i.e. conditions under which
$\Phi_{\Delta,\Delta^{\prime}}(\mathcal{M},\mathcal{M}^{\prime};\Omega,\Omega^{\prime})=\Phi_{\Delta}({\mathcal{M}};\Omega)+\Phi_{\Delta^{\prime}}({\mathcal{M}^{\prime}};\Omega^{\prime})$
holds. If so, we say that messages $M$ and $M^{\prime}$ are pragmatically independent with respect to their respective decision makers, $\Delta$ and $\Delta^{\prime}$, with the understanding that $\Delta$ and $\Delta^{\prime}$ could coincide. The following corollary to Theorem 3.2.3 gives a sufficient, but not necessary condition for such pragmatic independence:
Corollary (Additivity of Pragmatic Information).
If, in addition, to the hypotheses of Theorem 3.2.3, we have
[…]
then
$\Phi_{\Delta,\Delta^{\prime}}(\mathcal{M},\mathcal{M}^{\prime};\Omega,\Omega^{\prime})=\Phi_{\Delta}({\mathcal{M}};\Omega)+\Phi_{\Delta^{\prime}}({\mathcal{M}^{\prime}};\Omega^{\prime})$ (4)
Proof.
Summing over $i$ and $i^{\prime}$, we conclude that $\varphi_{m,m^{\prime}}=\varphi_{m}\varphi_{m^{\prime}}$. Thus,
$p_{i,i^{\prime}|m,m^{\prime}}$ $=\frac{p_{i,i^{\prime},m,m^{\prime}}}{\varphi_{m,m^{\prime}}}$
$=\left[\frac{p_{i,m}}{\varphi_{m}}\right]\left[\frac{p^{\prime}_{i^{\prime},m^{\prime}}}{\varphi_{m^{\prime}}}\right]$
$=p_{i|m}p^{\prime}_{i^{\prime}|m^{\prime}}.$
The result follows, since
$\Phi_{\Delta,\Delta^{\prime}}(\mathcal{M},\mathcal{M}^{\prime};\Omega,\Omega^{\prime})$ $=\sum_{i,i^{\prime},m,m^{\prime}}p_{i,i^{\prime},m,m^{\prime}}\log\left(\frac{p_{i,i^{\prime}|m,m^{\prime}}}{q_{i,i^{\prime}}}\right)$
$=\sum_{i,i^{\prime},m,m^{\prime}}p_{i,m}p^{\prime}_{i^{\prime},m^{\prime}}\log\left(\frac{p_{i|m}p_{i^{\prime}|m^{\prime}}}{q_{i}q_{i^{\prime}}}\right)$
$=\Phi_{\Delta}({\mathcal{M}};\Omega)+\Phi_{\Delta^{\prime}}({\mathcal{M}^{\prime}};\Omega^{\prime}).$
(3) is not necessary, because any message $m$ is pragmatically independent of any other message $m^{\prime}\in\mathcal{m}$ for which $D_{\Delta^{\prime}}(\mathbf{p}_{m^{\prime}}||\mathbf{q})=0$ for all $m^{\prime}\in\mathcal{M}^{\prime}$.
Nor is mere probabalistic independence of $m$ and $m^{\prime}$ sufficient to guarantee pragmatic independence, as can be seen from the example in the previous subsection. There, we see that
$\Phi_{\Delta,\Delta^{\prime}}(\mathcal{M},\mathcal{M}^{\prime};\Omega,\Omega^{\prime})\neq\Phi_{\Delta}({\mathcal{M}};\Omega)+\Phi_{\Delta^{\prime}}({\mathcal{M}^{\prime}};\Omega^{\prime})$
even though $M$ and $M^{\prime}$ of that subsection are probabilistically independent.
4 Pragmatic Information as “Free Information,” By Way of Mutual Information
Pragmatic information can be interpreted as a kind of information analog of the Hemholtz free energy of statistical mechanics in the following sense: the Hemholtz free energy, $F$, is given by the formula
$F=U-TS.$
It is is a measure of the amount of energy in a system available to do useful work at a constant temperature, $T$, as opposed to the total energy, $U$, of the system ($S$ is the system’s entropy). Similarly, as we now show, pragmatic information is what remains when information irrelevant to making the decision is removed.
Denote the mutual information between $\mathcal{M}$ and $\Omega$ as $\mathcal{I}(\mathcal{M};\Omega)$, $p_{i}=\sum_{m\in\mathcal{M}}p_{i,m}$ for all $i$, and $\mathbf{p}$ as the vector of these marginal probabilities [^5] Note the distinction between $\mathbf{p}_{m}$ and $\mathbf{p}$: the former is the probability distribution of $\omega$, given a specific $m$, but the latter is the average of these distributions over all $m$’s.. We then have the following
Theorem 4.0.1.
$\Phi_{\Delta}(\mathcal{M};\Omega)$ $=\mathcal{I}(\mathcal{M};\Omega)+D_{\Delta}(\mathbf{p}||\mathbf{q})$
$=\mathcal{H}(\mathcal{M})-H(\mathcal{M}|\Omega)+D_{\Delta}(\mathbf{p}||\mathbf{q})$
Proof.
$\Phi_{\Delta}(\mathcal{M};\Omega)$ $=\sum_{i,m}\varphi_{m}p_{i|m}\log\left[\frac{p_{i|m}}{q_{i}}\right]$
$=\sum_{i,m}p_{i,m}\log\left[\frac{p_{i|m}}{p_{i}}\frac{p_{i}}{q_{i}}\right]$
$=\sum_{i,m}p_{i,m}\log\left[\frac{p_{i|m}}{p_{i}}\right]+\sum_{i}p_{i}\log\left(\frac{p_{i}}{q_{i}}\right)$
$=\mathcal{I}(\mathcal{M};\Omega)+D_{\Delta}(\mathbf{p}||\mathbf{q}),$
$=\mathcal{H}(\mathcal{M})-H(\mathcal{M}|\Omega)+D_{\Delta}(\mathbf{p}||\mathbf{q})$
since $\sum_{i}p_{i}\log\left(\frac{p_{i}}{q_{i}}\right)$ is the Kullbach-Leibler Divergence of $\mathbf{p}$ and $\mathbf{q}$. ∎
We also have the immediate
Corollary.
$\Phi(\mathcal{M};\Omega)\geq\mathcal{I}(\mathcal{M};\Omega),$ with equality if and only if $\mathbf{p}=\mathbf{q}$ [^6] This corollary resolves a confusion in [3] by showing that pragmatic information and mutual information are equal only in this special case..
When $\mathbf{p}=\mathbf{q}$, the information needed to specify outcome $\omega_{i}\in\Omega$, i.e. the amount of information used in making a presumably useful decision, is the total amount of information transmitted by receipt of $m\in\mathcal{M}$, reduced by $H(\mathcal{M}|\Omega)$, the amount of information that is unneeded to specify the given outcome, and thus useless for making the decision. This difference is similar in spirit to free energy, the difference between the total energy and the internal energy, i.e. the information that is not available to do useful work. We can therefore characterize pragmatic information as “free information”; in fact, it is a generalization of the free information defined in [1].
In the more general case in which $\mathbf{p}\neq\mathbf{q}$, the term $D_{\Delta}(\mathbf{p}||\mathbf{q})$ is needed to update the biases that $\Delta$ had prior to the receipt of $m\in\mathcal{M}$. In this case, $\Delta$ is provided with two kinds of information; in addition to $\mathcal{H}(\mathcal{M})$, $\Delta$ is also provided with the information of the update. The characterization of pragmatic information as free information, i.e. the difference between total information and “irrelevant information”, still makes sense.
5 Pragmatic Utility and the Efficient Market Hypothesis
As indicated in the Introduction, a given message/message ensemble can, for a given decision maker, $\Delta$, have positive pragmatic information, yet misinform $\Delta$. This follows from our definition of pragmatic information as information that changes $\Delta$’s views of the state of the world, regardless of whether the new view is any more accurate, better, etc. than some base state. Nor is $\Delta$ guaranteed to have enough context to assign a value to that decision and thus to the messages leading to that decision. For example, what should be the value of a message that can prompt a single, terrible, but low probability outcome, as opposed to a message that leads to a set of more likely, but merely bad outcomes? Or can we be sure that we can assign values $V(m_{1}),V(m_{2}),V(m_{3})$ to messages $m_{1},m_{2},m_{3}\in\mathcal{M}$, such that $V(m_{1})<V(m_{2})$ and $V(m_{2})<V(m_{3})$ implies $V(m_{1})<V(m_{3})$?
Instead, we make the weaker assumption that we know enough about the implications of messages to classify them as being either
  • pragmatically irrelevant with respect to $\Delta$, or
  • pragmatic disinformation (or simply “disinformation”) with respect to $\Delta$,
  • pragmatically useful with respect to $\Delta$.
We might, for example, make the above classification based on whether a given message is less useful, the same as, or more useful to $\Delta$ than not receiving the message at all. In any case, (2) is the same for pragmatic disinformation and pragmatically useful information, because that formula captures only the degree to which the probabilties of $\Delta$’s decisions, and not the value of these decisions to $\Delta$, have changed.
For a given $\Delta$, we can then partition $\mathcal{M}$ as $\mathcal{M}=\mathcal{I}_{\Delta}\cup\mathcal{D}_{\Delta}\cup\mathcal{U}_{\Delta}$, where:
  • $\mathcal{I}_{\Delta}$ is the set of pragmatically irrelevant messages with respect to $\Delta$.
  • $\mathcal{D}_{\Delta}$ is the set of pagmatically disinformative messages with respect to $\Delta$.
  • $\mathcal{U}_{\Delta}$ is the set of pragmatically useful messages with respect to $\Delta$.
The corresponding amounts of pragmatically irrelevant, disinformative, and useful messages can then be defined, as $\Phi_{\Delta}(\mathcal{I}_{\Delta};\Omega),\Phi_{\Delta}(\mathcal{D}_{\Delta};\Omega)$, and $\Phi_{\Delta}(\mathcal{U}_{\Delta};\Omega)$, respectively, with
$\Phi_{\Delta}(\mathcal{I}_{\Delta};\Omega)=\sum_{m\in\mathcal{I}_{\Delta}}\sum_{i}p_{i,m}\log\left(\frac{p_{i|m}}{q_{i}}\right),$
$\Phi_{\Delta}(\mathcal{D}_{\Delta};\Omega)=\sum_{m\in\mathcal{D}_{\Delta}}\sum_{i}p_{i,m}\log\left(\frac{p_{i|m}}{q_{i}}\right),$
and
$\Phi_{\Delta}(\mathcal{U}_{\Delta};\Omega)=\sum_{m\in\mathcal{U}_{\Delta}}\sum_{i}p_{i,m}\log\left(\frac{p_{i|m}}{q_{i}}\right).$
Clearly,
$\Phi_{\Delta}(\mathcal{I}_{\Delta};\Omega),\Phi_{\Delta}(\mathcal{D}_{\Delta};\Omega),\Phi_{\Delta}(\mathcal{U}_{\Delta};\Omega)\geq 0,$
and
$\Phi_{\Delta}({\mathcal{M}};\Omega)=\Phi_{\Delta}(\mathcal{I};\Omega)+\Phi_{\Delta}(\mathcal{D};\Omega)+\Phi_{\Delta}(\mathcal{U};\Omega).$
It follows that
$\Phi_{\Delta}(\mathcal{I};\Omega),\Phi_{\Delta}(\mathcal{D};\Omega),\Phi_{\Delta}(\mathcal{U};\Omega)\leq\Phi_{\Delta}(\mathcal{M};\Omega).$
In particular, if the pragmatic information in some message ensemble $\mathcal{E}$ is zero, then it cannot contain any pragmatically useful information.
We now apply the above definitions to make a useful reformulation of the efficient market hypothesis, a cornerstone of financial economics (See, for example, [18].). According to this hypothesis, current asset prices, such as stock and bond prices, reflect all available information, so there is nothing to be gained by poring over charts of price histories or over annual reports.
However, the assumption that market efficiency is a property of asset price fluctuations suggests that all investors should find the markets equally inscrutable, a claim belied by the conspicuous long term success of such investors as Warren Buffett. If we take the pragmatic utility of market information to be its ability to improve its recipient’s ability to achieve above-market returns, we can therefore propose the following reformulation:
Hypothesis.
None of the information available to a given market participant is pragmatically useful.
Clearly, the above characterization of market efficiency depends on the participant, as well as the message ensemble being considered. The market can therefore be efficient to some participants, but not others because information may be pragmatically useful to one market participant but not others [^7] This is certainly the case, for example, regarding knowledge of short term mispricings that can be exploited by automated trading systems, but not human traders.. That the market can be efficient with respect to certain message ensembles, but not others is already a staple of the relevant literature. [18], for example, distinguishes between the “weak”, “semi-strong”, and “strong” forms of the hypothesis that, respectively, price histories, all available public information, and all available information, including insider information, are irrelevant in predicting future asset price fluctuations. Nevertheless, our conceptual framework seems to provide one immediate insight: namely that a market is necessarily efficient with respect to a given participant and a given message ensemble if that participant is unable to extract any pragmatically useful information from that ensemble. One such situation is when the amount of incoming market information overwhelms the computational capacities of the participant, as the present author has found from personal experience!
6 Directions for Future Research: $\Delta$’s Computational Capacity and Beyond
Implicit in the above are the limits that a decision maker’s computational abilities play in accruing pragmatic information. A sense of the importance of the role of these limits is made clear by an example due to [9]: If $\Delta$ has no computational abilities at all, in the sense that any message received will be ignored, $\mathbf{p}_{m}=\mathbf{q}$ for all messages $m$, and $\Phi_{\Delta}({\mathcal{M}};\Omega)=0$, regardless of the length or complexity of $m$! In other words, no computation implies no pragmatic information!
Furthermore, the details of the computational apparatus that $\Delta$ is assumed to possess are relevant, not just for our efficient markets application, but also for biological evolution. Recall, per [3] and references therein, that the quasi-species is a simplfied model of abstract “replicators”, subject to Darwinian “survival of the fittest” in the presence of mutation. It is the interaction of the quasi-species with its environment that “decides” which replicator is, in fact, the fittest. Therefore, this interaction is the $\Delta$ of the present theory, because it is this interaction that “decides” the unknown identity of the fastest growing replicator. The identity of this replicator is the random variable, $\omega$, and the environmental details that determine growth and mutation rates are the “messages” that the quasi-species environment system processes. The whole point of the formulation of the quasi-species theory [19] was to demonstrate the existence of an “error catastrophe”, which is a limit to the information that the quasi-species “computation” could reliably accrue in any given situation.
Also, consider the example of the one armed bandit. The above calculation assumes that we have the arbitrarily large amount of memory in which we can store the entire history of play. The calculation would look quite different if we stored, for example, only the most recent 5 trials.
All of the above suggests the need for a systematic exploration of how $\Delta$’s ability to accrue pragmatic information depends on its computational abilities. One approach might be to formalize these abilities via the Chomsky Hierarchy familiar to students of abstract automata [20] or the more detailed hierarchy in [1]. The latter author also suggests that the extensive semigroup theory of automata and his statistical mechanical approach might be of use in understanding how $\Delta$ partitions its input messages into its possible updated “state of the world” estimates.
In the present paper, we have assumed that a single message updates $\Delta$’s estimates of a single, discrete valued random variable $\omega$, the simplest possible setting for our ideas. For some applications, however, the assumption that $\omega$ is a continuous random variable, or even a stochastic process, is more appropriate. It might also be useful to assume that $\Delta$ is responding to a stream of messages by making multiple decisions, perhaps with the goal of rendering each successsive message more pragmatically useful. We might even imagine that $\Delta$’s computational machinery is, itself, being updated by these messages, whether via software updates, neural plasticity, or some other mechanism. Such generalizations might provide useful insights into the ubiquitous phenomenon of collective intelligence, of which efficient markets are but one example [^8] In fact, the recent Santa Fe Institute Conference on Collective Intelligence inspired this work. See the conference website at https://www.santafe.edu/info/collective-intelligence-2023/about/..
Acknowledgements
The author would like to acknowledge Profs. Harald Atmanspacher and Herbert Scheingraber for organizing the NATO Advanced Study Institute on Information Dynamics, where the author first learned of the need for a theory of pragmatic information; New York University for a partial stipend to attend the Santa Fe Institute Conference on Collective Intelligence that encouraged him to complete this work; Prof. Atmanspacher, Kevin Atteson, and Keith Lewis for useful discussions; and Wayne Mareci for proofreading the manuscript.